Procedures for LEM Deacons

SU (Set Up):

Arrive at 4:45 PM. Supplies are downstairs in the two small rooms used by LEM; keys are on hooks in the grooves beside the doors to these rooms. Set up the alter with the appropriate alter cloth (refer to the roster for the color for the week). On top of it, place the white cloth and 2 communion cloths. Choose a cross and two candles to go on top or in front of the alter. Light the candles (matches are available downstairs). Bring up bread and wine, put them at the back on the ledge. Put the “Service in Progress” sign by the chapel entrance. Set up the table with service books, hymnals, namecards and pen (refer to the roster for which service books and hymnals to use); these are kept downstairs.

After the service, put everything away and wash the communion ware.

RC (Readers and Communion):

Our Lutheran vicar, Margot, will provide you with the readings several days before the service so that you can read them over and become familiar with them.

On the day of the service, arrive at 5 PM to obtain the lectionary and place bookmarks at the locations of the readings. Before the service starts, find someone who will help you bring forward the bread and wine for Communion when the time comes later in the service.

During the service, conclude each reading with “The Word of God” or “The Word of the Lord”. Serve as Communion Assistant by serving one of the two wine chalices. When administering the wine to a recipient, state “This is the Blood of Christ, shed for you” or some other appropriate phrase.

After the service, please help the setup person clean up after the service.

PP (Prayers of the People):

Write 4 to 7 intercessory prayers. Say them at the appropriate time. Conclude each intercession with “God in your mercy,” (congregation answers “Hear our prayer”) or “Bless we our God” (congregation answers “Thanks be to God”). In the first concluding phrase, you may wish to substitute “Gracious Mother”, “Gracious Father” or “Lord” for the word “God”. In the second case, you might prefer to use “the Lord” instead of “our God”. Between each prayer, leave a few moments of silence when people can medidate and offer their own related prayers.

Absences:

Please try to be present on the days when you are scheduled to serve. If you have a schedule conflict and cannot attend the service, please try to find someone who will substitute for you; you can also try phoning John Keen (225-9841) or Margot Vander Ziel (876-4996) and one of us will help you out.

Further Questions?

If you need any clarifications or elaborations on any of these procedures, please ask John, Margot, Scott or Susan and we will be happy to answer any questions you may have.

Thank you for serving the Lord’s community in the Lutheran-Episcopal Ministry at MIT!